OpenStudy (anonymous):

if a third term of an arithmetic is 14 and the 8th term is -1, find the values of a1 and d of the sequence

OpenStudy (anonymous):

How many steps from the 3rd term and the 8th term?

OpenStudy (anonymous):

d=-5/2

OpenStudy (anonymous):

a1 is 19

OpenStudy (anonymous):

how do you do it

OpenStudy (anonymous):

Just write the series dude a1, a1+d, a1+2d(3rd term=14) , a1+3d, ........a1+7d(8th term=-1) Solve two equations for a1 and d a1+2d=14 a1+7d=-1 Is it ok?????

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

Got it dude????????

OpenStudy (anonymous):

not yet they do not want me to list them

OpenStudy (anonymous):

Didnt got u

OpenStudy (anonymous):

U have been given with the 3rd and 8th terms to find first term and arith.difference of the series right????? Then use the formula u have for nth term a+(n-1)d then substitute n value for ur terms given and eqaute to their vale Solve the equations and get the result This will the same thing i have done above

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

sorry but I do not have a1 and d. and whats n

OpenStudy (anonymous):

n is term given to u for the third term it is 3 and 8th term it is 8 and the expression in each equals to the value given to you

OpenStudy (anonymous):

is it third=14 and 8th=-1

OpenStudy (anonymous):

yeah exactly a+(3-1)d=14 a+(8-1)d=-1

OpenStudy (anonymous):

thats the method of two equation then sub 2 into 1 am I right

OpenStudy (anonymous):

yeah i think you know to solve two simultaneous equations right????

OpenStudy (anonymous):

yes I just did not get it at the first. I will let you know how I went

OpenStudy (anonymous):

It appears like the delta between terms is -3. Assuming a1 is the first term of the series, it has to be 14+3+3 = 20.

OpenStudy (anonymous):

I think I did it d=3 and a1=8

OpenStudy (anonymous):

Atchyut are you there?

OpenStudy (anonymous):

how does this work I did the second time and got d=5/3=1.6 and a1=10.8. both work so which one to use. help please

OpenStudy (anonymous):

ok wait i will tell

OpenStudy (anonymous):

you have the two equations a1+2d=14 a1+7d=-1 right????

OpenStudy (anonymous):

yes did and the results are the ones written above

OpenStudy (anonymous):

Subtract the second equation from first then a1 and a1 gets cancelled we will remain with 2d-7d=14-(-1) -5d=15 and d= -15/5 =-3 and now substitute d=-3 in first equation a1-6=14 gives a1=20. So a1 is 20 and d=-3 At first I thought something wrong and kept pth term equation inplace of 8 th term and so we get it like that

OpenStudy (anonymous):

ok I will do it again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how did you get -3 instead of 3

OpenStudy (anonymous):

Once check my solution dude

OpenStudy (anonymous):

think once first term is greater than the next terms.. how this is possible????? It is possible only when the common difference is negative dude got it????

OpenStudy (anonymous):

its geting more confusing because I did and got d=3 and a1=8 and tryed to do the sequence and it works but does not work for 8th term=-1